OpenStudy (anonymous):

31+49 cos(theta)=49 sin^2(Theta)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

..... /49 31/49 + cos(t) = sin^2(t) for a start?

OpenStudy (amistre64):

sin^2(t) = sqrt(1 - cos^2(t)) sooo.... 31/49 + cos(t) = sqrt(1 - cos^2(t))

OpenStudy (amistre64):

(31/49)^2 + cos^2(t) + (62/49)cos(t) = 1 - cos^2(t)

OpenStudy (amistre64):

2cos^2(t) + (62/49)cos(t) +1 + (31/49)^2 = 0

OpenStudy (amistre64):

u = cos(t) 2u^2 + (62/49)u +(1 + (31/49)^2) = 0

OpenStudy (anonymous):

where did the 32/49 come from lol, there wasnt even a fraction to misinterperate!

OpenStudy (amistre64):

divide it all by 49 in the begining

OpenStudy (amistre64):

2u^2 + (62/49)u +(-1 + (31/49)^2) = 0 typoed lol